OpenStudy (anonymous):

multiply and write the answer without exponents. (3*10^5)*(3*10^6)

jimthompson5910 (jim_thompson5910):

Multiply the numbers in front (3 and 3) to get 3*3 = 9

jimthompson5910 (jim_thompson5910):

Then add the exponents (5 and 6) to get 5+6=11

jimthompson5910 (jim_thompson5910):

So (3*10^5)*(3*10^6) = 9*10^11 Use a calculator to expand that out Optionally, you can just move the decimal point 11 spots to the right

OpenStudy (anonymous):

900000000000

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np